# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ Appendix A Pseudo-stationarity
where $L$ and $\mu$ are defined in 5.1 for the $h$ we selected. Next, applying 4.1, we know for all $i\geq r+1$, it holds that $d_{i}^{1}\leq\lambda+L\sigma_{r}(X_{1})$; in addition, it must hold that $\|d^{1}\|_{\infty}>\lambda$ for otherwise, B.2 and 2.1 would imply that $X_{1}$ is a stationary point of $f$, which is a contradiction. On the other hand, using the fact that $-\nabla h(X_{2})\in\lambda\partial\|X_{2}\|_{*}$, B.2 and 2.1, we know $d_{i}^{2}\leq\lambda$ for all $i\in[m]$. This means that $(x,g,y,v,\tau)=(\mathrm{diag}(\Sigma_{1}),\mathrm{diag}(D_{1}),\mathrm{diag}(\Sigma_{2}),\mathrm{diag}(D_{2}),\bar{\tau})$ satisfies (B.1a)–B.1c and $\|g\|_{\infty}>\lambda$. Therefore we know that (ii) holds.
Next, assume (ii). Then, we are able to select $(x,g,y,v,\tau)$ satisfying B.1aB.1c. We sort $x$ and $y$ in descending order to get $\bar{x}$ and $\bar{y}$, respectively. Pick $\tau_{1},\tau_{2}\in\mathfrak{P}_{m}$ such that
$\bar{x}=\tau_{1}(x),~{}\bar{y}=\tau_{2}(y)\text{ and we define }\bar{g}=\tau_{1}(g),~{}\bar{v}=\tau_{2}(v).$ (B.4)
Since we have in view of B.1b and B.1c that
$\forall i\in[r],~{}x_{i}>0,~{}\quad\forall i\in[m]\setminus[r],~{}x_{i}=0,$ (B.5)
$\forall i\in[r^{*}],~{}y_{i}>0,~{}\quad\forall i\in[m]\setminus[r^{*}],~{}y_{i}=0,$
we can see from B.4 that
$\tau_{1}([r])=[r],\quad\tau_{1}([m]\setminus[r])=[m]\setminus[r],\quad\tau_{2}([r^{*}])=[r^{*}],\quad\tau_{2}([m]\setminus[r^{*}])=[m]\setminus[r^{*}].$ (B.6)
Moreover, notice that from B.1b and B.1c we have
$\forall i\in[r],~{}g_{i}=\lambda,~{}\quad\forall i\in[m]\setminus[r],~{}g_{i}\in[0,\lambda+L\min_{i\in[r]}x_{i}],$ (B.7)
$\forall i\in[r^{*}],~{}v_{i}=\lambda,~{}\quad\forall i\in[m]\setminus[r^{*}],~{}v_{i}\in[0,\lambda],$
This together with the definitions of $\bar{g}$ and $\bar{v}$ in B.4 and the relations B.5 and B.6 implies that
$\forall i\in[r],~{}\bar{x}_{i}>0,~{}\bar{g}_{i}=\lambda,~{}\quad\forall i\in[m]\setminus[r],~{}\bar{x}_{i}=0,\bar{g}_{i}\in[0,\lambda+L\min_{i\in[r]}\bar{x}_{i}],$ (B.8)
$\forall i\in[r^{*}],~{}\bar{y}_{i}>0,~{}\bar{v}_{i}=\lambda,~{}\quad\forall i\in[m]\setminus[r^{*}],~{}\bar{y}_{i}=0,~{}\bar{v}_{i}\in[0,\lambda].$
Let $\rho\in\mathfrak{P}_{m}$ be defined as $\rho:=\tau_{2}^{-1}\tau\tau_{1}$ and define the following $m\times n$ matrices:
$X_{1}=\widetilde{\rm Diag}(\bar{x}),~{}X_{2}=\widetilde{\rm Diag}(\rho(\bar{y})),~{}G_{1}=\widetilde{\rm Diag}(\bar{g}),~{}G_{2}=\widetilde{\rm Diag}(\rho(\bar{v})).$ (B.9)
By direct calculation, we have:
$\sum_{i=1}^{m}(L\sigma_{i}(X_{1})+\bar{g}_{i})(\mu\sigma_{\rho(i)}(X_{2})+\bar{v}_{\rho(i)})+\sum_{i=1}^{m}(\mu\sigma_{i}(X_{1})+\bar{g}_{i})(L\sigma_{\rho(i)}(X_{2})+\bar{v}_{\rho(i)})$
$-\sum_{i=1}^{m}(L\sigma_{i}(X_{1})+\bar{g}_{i})(\mu\sigma_{i}(X_{1})+\bar{g}_{i})-\sum_{i=1}^{m}(L\sigma_{i}(X_{2})+\bar{v}_{i})(\mu\sigma_{i}(X_{2})+\bar{v}_{i})$
=$\overset{\rm(a)}{=}$ $\sum_{i=1}^{m}(L\bar{x}_{i}+\bar{g}_{i})(\mu\bar{y}_{\rho(i)}+\bar{v}_{\rho(i)})+\sum_{i=1}^{m}(\mu\bar{x}_{i}+\bar{g}_{i})(L\bar{y}_{\rho(i)}+\bar{v}_{\rho(i)})$
$-\sum_{i=1}^{m}(L\bar{x}_{i}+\bar{g}_{i})(\mu\bar{x}_{i}+\bar{g}_{i})-\sum_{i=1}^{m}(L\bar{y}_{i}+\bar{v}_{i})(\mu\bar{y}_{i}+\bar{v}_{i})$
=$\overset{\rm(b)}{=}$ $\sum_{i=1}^{m}(Lx_{\tau_{1}(i)}+g_{\tau_{1}(i)})(\mu y_{\tau_{2}(\rho(i))}+v_{\tau_{2}(\rho(i))})+\sum_{i=1}^{m}(\mu x_{\tau_{1}(i)}+g_{\tau_{1}(i)})(Ly_{\tau_{2}(\rho(i))}+v_{\tau_{2}(\rho(i))})$
$-\sum_{i=1}^{m}(Lx_{\tau_{1}(i)}+g_{\tau_{1}(i)})(\mu x_{\tau_{1}(i)}+g_{\tau_{1}(i)})-\sum_{i=1}^{m}(Ly_{\tau_{2}(i)}+v_{\tau_{2}(i)})(\mu y_{\tau_{2}(i)}+v_{\tau_{2}(i)})$
=$\overset{\rm(c)}{=}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$
$-\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu x_{i}+g_{i})-\sum_{i=1}^{m}(Ly_{i}+v_{i})(\mu y_{i}+v_{i})\overset{\rm(d)}{\geq}0,$ (B.10)
where in (a) we have used the fact that $\bar{x}$ and $\bar{y}$ are nonnegative vectors sorted in descending order to calculate $\sigma_{i}(X_{1})$ and $\sigma_{i}(X_{2})$ using B.9, in (b) we have used B.4, in (c) we have used that $\rho=\tau_{2}^{-1}\tau\tau_{1}$, the substitution $i\leftarrow\tau_{1}^{-1}(i)$ for the first three terms, and the substitution $i\leftarrow\tau_{2}^{-1}(i)$ for the last term, and in (d) we have used B.1a. Applying Lemma A.3, and noticing that the prerequisites in Lemma A.3 are satisfied by B.8, Appendix B and the fact that $\bar{x}$ and $\bar{y}$ are nonnegative vectors sorted in descending order, we know there exists a quadratic $h\in C^{2}(\mathbb{R}^{m\times n})$, such that $\nabla h(X_{i})=-G_{i}$ for $i=1,2$, $h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$ is convex, and $\nabla h$ is Lipschitz continuous with modulus $L$. Moreover, $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$ in 1.5, where $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ is defined as:
@@
$\bar{U}=\widetilde{\rm Diag}(\sqrt{\sigma_{1}(X_{1})},\dots,\sqrt{\sigma_{r}(X_{1})}),\quad\bar{V}=\widetilde{\rm Diag}(\sqrt{\sigma_{1}(X_{1})},\dots,\sqrt{\sigma_{r}(X_{1})}).$ (B.11)
However, since $\|\bar{g}\|_{\infty}>\lambda$, we know $-\nabla h(X_{1})=G_{1}\notin\lambda\partial\|X_{1}\|_{*}$ by 2.1. This shows that $X_{1}$ is not a stationary point of $f$ in 1.4, and hence $f$ is not $r$-factorizable. Finally, in view of the definitions of $X_{2}$ and $G_{2}$ in B.9, the relations in B.8 and 2.1, we can deduce that $-\nabla h(X_{2})=G_{2}\in\lambda\partial\|X_{2}\|_{*}$ and $\mathrm{rank}(X_{2})=r^{*}$, from which we conclude further that $h\in\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$. Consequently, we know (i) holds. ∎
Let $\infty>L\geq\mu>0$. Consider the next optimization problem:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m}\end{subarray}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.12)

### v2

@@ 4 Characterization of $r$-factorizability for strongly convex functions
where $L$ and $\mu$ are defined in Definition 1.2 for the $h$ we selected. Next, applying 3.7, we know for all $i\geq r+1$, it holds that $d_{i}^{1}\leq\lambda+L\sigma_{r}(X_{1})$; in addition, it must hold that $\|d^{1}\|_{\infty}>\lambda$ for otherwise, 4.2 and Proposition 2.1 would imply that $X_{1}$ is a stationary point of $f$, which is a contradiction. On the other hand, using the fact that $\nabla h(X_{2})\in\lambda\partial\|X_{2}\|_{*}$, 4.2 and Proposition 2.1, we know $d_{i}^{2}\leq\lambda$ for all $i\in[m]$. This means that $(x,g,y,v,\tau)=(\mathrm{diag}(\Sigma_{1}),\mathrm{diag}(D_{1}),\mathrm{diag}(\Sigma_{2}),\mathrm{diag}(D_{2}),\bar{\tau})$ satisfies (4.1a)–4.1c and $\|g\|_{\infty}>\lambda$. ∎
To provide a complete characterization, it is important to know whether the converse of Proposition 4.1 holds. The first issue we need to solve is that, a feasible pair $(x,g,y,v)$ satisfying 4.1b and 4.1c does not necessarily satisfy that $x$ and $y$ are sorted in descending order, which means that they cannot be the singular values of any matrix. Nevertheless, we can reduce to the case where $(x,y)$ are sorted in descending order by doing an additional permutation.
Lemma 4.2.
Let $L\geq\mu>0$, $r\in[m]$ and $r^{*}\in[m]\cup\{0\}$. Assume that $(x,g,y,v,\tau)$ satisfies 4.1a4.1c with $\|g\|_{\infty}>\lambda$, and we pick $\tau_{1},\tau_{2}\in\mathfrak{P}_{m}$ such that [^8] Here for a vector $x\in\mathbb{R}^{n}$ and a permutation $\tau$, $\tau(x)\in\mathbb{R}^{n}$ is defined as $\tau(x)_{i}=x_{\tau(i)}$ for all $i\in[n]$.
[…]
Then $(\bar{x},\bar{g},\bar{y},\bar{v},\rho)$ also satisfies 4.1a4.1c with $\|\bar{g}\|_{\infty}>\lambda$, where $\rho:=\tau_{2}^{-1}\tau\tau_{1}$.
Proof.
According to the constraints on $x,y$ in 4.1b and 4.1c, we know that $x_{i}>0$ if and only if $i\in[r]$, and $y_{j}>0$ if and only if $j\in[r^{*}]$. Consequently, according to the definition of $\tau_{1}$ and $\tau_{2}$, we have
$\tau_{1}([r])=[r],\quad\tau_{1}([m]\setminus[r])=[m]\setminus[r],\quad\tau_{2}([r^{*}])=[r^{*}],\quad\tau_{2}([m]\setminus[r^{*}])=[m]\setminus[r^{*}].$ (4.5)
Therefore, $(\bar{x},\bar{g},\bar{y},\bar{v})$ satisfies 4.1b and 4.1c. To verify 4.1a, notice that the four sums in 4.1a takes the form $\sum_{i=1}^{m}a_{i}b_{\tau(i)}$ for some $a,b\in\mathbb{R}^{m}$. We also note that for any $a,b,c,d\in\mathbb{R}^{m}$ with $a=\tau_{1}(c)$ and $b=\tau_{2}(d)$, it holds that
$\sum_{i=1}^{m}a_{i}b_{\rho(i)}\overset{\rm(a)}{=}\sum_{i=1}^{m}c_{\tau_{1}(i)}d_{\tau_{2}\rho(i)}\overset{\rm(b)}{=}\sum_{i=1}^{m}c_{\tau_{1}(i)}d_{\tau\tau_{1}(i)}\overset{\rm(c)}{=}\sum_{i=1}^{m}c_{i}d_{\tau(i)},$ (4.6)
where (a) holds because $a=\tau_{1}(c)$ and $b=\tau_{2}(d)$, (b) holds as $\rho=\tau_{2}^{-1}\tau\tau_{1}$, and in (c) we have used the substitution $i\leftarrow\tau_{1}^{-1}(i)$. Consequently, to verify 4.1a for $(\bar{x},\bar{g},\bar{y},\bar{v},\rho)$, it suffices to use 4.4, 4.6 and 4.1a for $(x,g,y,v,\tau)$. Finally, the condition $\|\bar{g}\|_{\infty}>\lambda$ is clear since $\bar{g}$ is a permutation of $g$ and $\|g\|_{\infty}>\lambda$. ∎
Consequently, if there exists $(x,y,g,v,\tau)$ satisfying 4.1a4.1c with $\|g\|_{\infty}>\lambda$, then there is no loss of generality if we assume $x$ and $y$ are sorted in descending order. In this case, we would like to construct $\bar{X}$, $\nabla h(\bar{X})$, $X^{*}$ and $\nabla h(X^{*})$ such that $\sigma(\bar{X})=x$, $\sigma(X^{*})=y$ and the bound in 1.8 is satisfied. To this end, we first define
$\bar{X}=\widetilde{\rm Diag}(x),~X^{*}=\widetilde{\rm Diag}(\tau(y)),~\bar{G}=\widetilde{\rm Diag}(g),~G^{*}=\widetilde{\rm Diag}(\tau(v)),$ (4.7)
In this case, by direct calculation, we have
$(L-\mu)\langle(G^{*}+\mu X^{*})-(\bar{G}+\mu\bar{X}),\bar{X}-X^{*}\rangle-\|(\bar{G}+\mu\bar{X})-(G^{*}+\mu X^{*})\|_{F}^{2}$ (4.8)
$=\langle(G^{*}+\mu X^{*})-(\bar{G}+\mu\bar{X}),L\bar{X}+\bar{G}-(LX^{*}+G^{*})\rangle$
$=\langle G^{*}+\mu X^{*},L\bar{X}+\bar{G}\rangle+\langle\bar{G}+\mu\bar{X},LX^{*}+G^{*}\rangle$
$~~~~~-\langle G^{*}+\mu X^{*},LX^{*}+G^{*}\rangle-\langle\bar{G}+\mu\bar{X},L\bar{X}+\bar{G}\rangle$
[…]
$~~~~-\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu x_{i}+g_{i})-\sum_{i=1}^{m}(Ly_{i}+v_{i})(\mu y_{i}+v_{i})\overset{\rm(b)}{\geq}0$
where (a) follows from the diagonal structures in 4.7, and (b) follows from 4.1a. The display 4.8 implies that letting $\nabla h(\bar{X})=-\bar{G}$ and $\nabla h(X^{*})=-G^{*}$ does not violate 1.8 when $h$ is $\mu$-strongly convex and $\nabla h$ is $L$-Lipschitz. The next question is, does there exist such a function $h\in C^{2}(\mathbb{R}^{m\times n})$? Notice that $\nabla h(X^{*})\in\lambda\partial\|X^{*}\|_{*}$ if such a function exists, which means $X^{*}$ is a stationary point of $f$ in 1.1, and $h\in\mathfrak{S}(L,\mu,r^{*})$. This question is a special case of the extension of strongly convex function [1]. However, for our purpose, we need to further ensure that there exists a second-order stationary point $(\bar{U},\bar{V})$ of the corresponding $F_{r}$ in 1.2 such that $\bar{U}\bar{V}^{\top}=\bar{X}$, which will be achieved by adding quadratic penalty on the off diagonal entries due to the structure of the second-order optimality condition in Section 3. Let us now provide the construction of such example.
Lemma 4.3.
Let $L\geq\mu>0$, $r\in[m]$ and $r^{*}\in[m]\cup\{0\}$. Assume that $(x,g,y,v,\tau)$ satisfies 4.1a4.1c, and $x$ and $y$ are sorted in descending order. Let $\bar{X}$, $X^{*}$, $\bar{G}$, $G^{*}\in\mathbb{R}^{m\times n}$ be defined as 4.7. If $G^{*}+\mu X^{*}\neq\bar{G}+\mu\bar{X}$, we define a quadratic function $h$ as follows:
$h(X)=$ $\frac{L}{2}\sum_{i=1}^{m}\sum_{j\neq i}^{n}X_{ij}^{2}+\frac{\mu}{2}\sum_{i=1}^{m}(X_{ii}-(\bar{X})_{ii})^{2}-\langle\bar{G},X\rangle$ (4.9)
$+\frac{(\langle X-\bar{X},-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\rangle)^{2}}{2\langle X^{*}-\bar{X},-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\rangle},$
Otherwise, we set
$h(X)=$ $\frac{L}{2}\sum_{i=1}^{m}\sum_{j\neq i}^{n}X_{ij}^{2}+\frac{\mu}{2}\sum_{i=1}^{m}(X_{ii}-(\bar{X})_{ii})^{2}-\langle\bar{G},X\rangle.$ (4.10)
Then $h$ is well defined and $\mu$-strongly convex, $\nabla h$ is Lipschitz continuous with modulus $L$, $\nabla h(\bar{X})=-\bar{G}$, and $\nabla h(X^{*})=-G^{*}$. Moreover, if we define $F_{r}$ as in 1.2 with the above $h$ and define $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ as
$\bar{U}:=\widetilde{\rm Diag}(\sqrt{\sigma_{1}(\bar{X})},\dots,\sqrt{\sigma_{r}(\bar{X})})\ \ {and}\ \ \bar{V}:=\widetilde{\rm Diag}(\sqrt{\sigma_{1}(\bar{X})},\dots,\sqrt{\sigma_{r}(\bar{X})}),$ (4.11)
then $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$. In particular, if $\|g\|_{\infty}>\lambda$, then $f$ in 1.1 is not $r$-factorizable given such an $h$.
Proof.
We first consider the case where $\bar{G}+\mu\bar{X}=G^{*}+\mu X^{*}$. In this case, the function $h$ in 4.10 is clearly well defined, and one can verify that $h$ is $\mu$-strongly convex, and $\nabla h$ is Lipschitz continuous with modulus $L$. Moreover, $\nabla h(\bar{X})=-\bar{G}$ and
$\nabla h(X^{*})=\mu(X^{*}-\bar{X})-\bar{G}=-G^{*}.$
We start by noticing from Proposition 3.2 that $(\bar{U},\bar{V})$ is a stationary point of $F_{r}$ (with $R=I_{m}$, $Q=I_{r}$ and $P=I_{n}$ in Proposition 3.2). Consequently, by Proposition 3.6, we know that $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$ if and only if for all $U_{11},V_{11}\in\mathbb{R}^{r\times r},~U_{21}\in\mathbb{R}^{(m-r)\times r},~V_{21}\in\mathbb{R}^{(n-r)\times r}$, [^9] Here, we use the partition in (3.14), which is well defined because $\sigma(\bar{U})=\sigma(\bar{V})$. We also note that $U_{12},V_{12},U_{22},V_{22}$ are void because $\mathrm{rank}(\bar{U})=r$. it holds that
$-2\lambda\mathrm{tr}(U_{11}^{\top}V_{11})-2\mathrm{tr}(D^{\top}U_{21}V_{21}^{\top})+\lambda(\|U_{11}\|_{F}^{2}+\|V_{11}\|_{F}^{2}+\|U_{21}\|_{F}^{2}+\|V_{21}\|_{F}^{2})$ (4.12)
$+\nabla^{2}h(\bar{X})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}\geq 0,$
where
@@
$\Sigma=\mathrm{diag}(\sqrt{\sigma_{1}(\bar{X})},\dots,\sqrt{\sigma_{r}(\bar{X})})\overset{(a)}{=}\mathrm{diag}(\sqrt{x_{1}},\dots,\sqrt{x_{r}})\in\mathbb{R}^{r\times r},$ (4.13)
$D=\widetilde{\rm Diag}(g_{r+1},\dots,g_{m})\in\mathbb{R}^{(m-r)\times(n-r)},$
and in (a) we have used the fact that $x$ is a nonnegative vector sorted in descending order. We will verify (4.12).
To this end, we first use the representation of $h$ in 4.10 to deduce that
$\begin{split}\nabla^{2}h(\bar{X})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}&\geq L\|\Sigma V_{21}^{\top}\|_{F}^{2}+L\|U_{21}\Sigma\|_{F}^{2}\\ &\overset{\rm{(a)}}{\geq}Lx_{r}(\|U_{21}\|^{2}_{F}+\|V_{21}\|_{F}^{2}),\end{split}$ (4.14)
where in (a) we have used the fact that $\|AB\|_{F}\geq\sigma_{\min}(A)\|B\|_{F}$ for any square matrix $A$. Therefore, it holds that
$+\nabla^{2}h(\bar{X})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}$
$\overset{\rm{(a)}}{\geq}-2\mathrm{tr}(D^{\top}U_{21}V_{21}^{\top})+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+\nabla^{2}h(\bar{X})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}$
$\overset{\rm{(b)}}{\geq}-2\|D^{\top}\|_{2}\|U_{21}\|_{F}\|V_{21}^{\top}\|_{F}+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+\nabla^{2}h(\bar{X})\left[\begin{bmatrix}U_{11}\Sigma+\Sigma V_{11}^{\top}&\Sigma V_{21}^{\top}\\ U_{21}\Sigma&0\end{bmatrix}\right]^{2}$
$\overset{\rm{(c)}}{\geq}-2(\lambda+Lx_{r})\|U_{21}\|_{F}\|V_{21}\|_{F}+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+Lx_{r}(\|U_{21}\|^{2}_{F}+\|V_{21}\|_{F}^{2})$
$\overset{\rm{(d)}}{\geq}-(\lambda+Lx_{r})(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+\lambda(\|U_{21}\|_{F}^{2}+\|V_{21}\|^{2}_{F})+Lx_{r}(\|U_{21}\|^{2}_{F}+\|V_{21}\|_{F}^{2})=0,$
where in (a) we have used the Cauchy-Schwartz inequality to show that $\mathrm{tr}(U_{11}^{\top}V_{11})\leq\frac{1}{2}(\|U_{11}\|^{2}_{F}+\|V_{11}\|_{F}^{2})$, in (b) we have used the fact $\mathrm{tr}(ABC)=\mathrm{tr}(CAB)\leq\|C\|_{F}\|AB\|_{F}\leq\|A\|_{2}\|C\|_{F}\|B\|_{F}$, in (c) we have used 4.1c, 4.14 and the definition of $D$ in 4.13, and in (d) we have used the fact $\|A\|_{F}\|B\|_{F}\leq\frac{1}{2}(\|A\|_{F}^{2}+\|B\|_{F}^{2})$. This verifies (4.12) and hence $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$.
Next, we consider the case where $\bar{G}+\mu\bar{X}\neq G^{*}+\mu X^{*}$. By 4.8 and the fact that $\bar{G}+\mu\bar{X}\neq G^{*}+\mu X^{*}$, we see that $(L-\mu)\langle(G^{*}+\mu X^{*})-(\bar{G}+\mu\bar{X}),\bar{X}-X^{*}\rangle\geq\|(\bar{G}+\mu\bar{X})-(G^{*}+\mu X^{*})\|_{F}^{2}>0$. In particular, this implies $L>\mu$ and $\langle(G^{*}+\mu X^{*})-(\bar{G}+\mu\bar{X}),\bar{X}-X^{*}\rangle>0$, showing that $h$ in 4.9 is well defined. Furthermore, we have
$L-\mu\geq\frac{\|(\bar{G}+\mu\bar{X})-(G^{*}+\mu X^{*})\|_{F}^{2}}{\langle(G^{*}+\mu X^{*})-(\bar{G}+\mu\bar{X}),\bar{X}-X^{*}\rangle}.$ (4.15)
Now, it is routine to check that $h$ is $\mu$-strongly convex, $\nabla h(\bar{X})=-\bar{G}$, and $\nabla h(X^{*})=-G^{*}$. Moreover, the relation in (4.14) and hence the second-order stationarity of $(\bar{U},\bar{V})$ can be verified similarly to that in the case where $\bar{G}+\mu\bar{X}=G^{*}+\mu X^{*}$. Thus, it remains to show that $\nabla h$ is Lipschitz continuous with modulus $L$.
To this end, notice that for all $X,Y\in\mathbb{R}^{m\times n}$ and the function $h$ defined in 4.9, it holds that
$\nabla^{2}h(X)[Y,Y]=L\sum_{i=1}^{m}\sum_{j\neq i}^{n}Y_{ij}^{2}+\mu\sum_{i=1}^{n}Y_{ii}^{2}+\frac{(\langle Y,-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\rangle)^{2}}{\langle X^{*}-\bar{X},-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\rangle}$
$\overset{\rm{(a)}}{=}L\sum_{i=1}^{m}\sum_{j\neq i}^{n}Y_{ij}^{2}+\mu\sum_{i=1}^{n}Y_{ii}^{2}+\frac{(\langle\widetilde{\rm Diag}(Y_{11},\dots,Y_{mm}),-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\rangle)^{2}}{\langle X^{*}-\bar{X},-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\rangle}$
$\overset{\rm(b)}{\leq}L\sum_{i=1}^{m}\sum_{j\neq i}^{n}Y_{ij}^{2}+\mu\sum_{i=1}^{n}Y_{ii}^{2}+\frac{\|\widetilde{\rm Diag}(Y_{11},\dots,Y_{mm})\|^{2}_{F}\,\|-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\|_{F}^{2}}{\langle X^{*}-\bar{X},-G^{*}-\mu X^{*}+\bar{G}+\mu\bar{X}\rangle}\overset{\rm(c)}{\leq}L\|Y\|_{F}^{2},$
where in (a) we have used the fact that $\bar{X},X^{*},\bar{G},G^{*}$ are diagonal (see 4.7), in (b) we have used the Cauchy-Schwartz inequality, and in (c) we have used (4.15). This proves that $\nabla h$ is Lipschitz continuous with modulus $L$.
Finally, if $\|g\|_{\infty}>\lambda$, then by Proposition 2.1 we know $\bar{X}$ is not a stationary point of $f$ in 1.1 given such an $h$, and hence cannot be a global minimizer. Therefore, in this case, $f$ in 1.1 is not $r$-factorizable. ∎
Combining Proposition 4.1, Lemma 4.2 and Lemma 4.3, we have the following result concerning the existence of the function $f$ in 1.1 when $h\in\mathfrak{S}(L,\mu,r^{*})$, which is not $r$-factorizable.
Proposition 4.4.
Let $L\geq\mu>0$, $r\in[m]$ and $r^{*}\in[m]\cup\{0\}$. Consider the following optimization problem:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m}\end{subarray}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (4.16)